BAB is a right isosceles triangle at D
1)Place point C so that DABC would be a parallelogram
2)The bisector of angle ADB cots AB at E and the bisector of angle DBC cuts CD at F
Prove that quadrilateral DEBF is a square
1
Expert's answer
2020-10-26T16:54:54-0400


Given: ∆DAB is a right isosceles triangle at D.

To prove: quadrilateral DBEF is a square.

Proof:

With the lettering in the diagram above,

"A\widehat{D}B=90°(\textsf{∆DAB is a right isosceles ∆ at D})\\
D\widehat{A}B=D\widehat{B}A(\textsf{base angles of an isosceles ∆})\\
D\widehat{A}B+D\widehat{B}A+A\widehat{D}B=180°\\(\textsf{angle sum of a ∆)}\\
D\widehat{A}B+D\widehat{B}A=180°-A\widehat{D}B\\
D\widehat{A}B+D\widehat{B}A=180°-90°\\
D\widehat{A}B+D\widehat{B}A=90°\\
D\widehat{A}B+D\widehat{A}B=90°(D\widehat{A}B=D\widehat{B}A)\\
2D\widehat{A}B=90°\\
D\widehat{A}B=45°\\" Since C is placed at a point so that DABC is a parallelogram, ∆BCD has to be a right isosceles ∆ which is similar to ∆DAB. Hence,

"D\widehat{B}C=A\widehat{D}B=90°\\
B\widehat{C}D=B\widehat{D}C=45°\\"

Since line DE bisects ADB,

"B\widehat{D}E=45°\\
E\widehat{D}F=B\widehat{D}C+B\widehat{D}E=45°+45°=90°"

Also, since line BF bisects "D\widehat{B}C,"

"D\widehat{B}F=45°\\
E\widehat{B}F=D\widehat{B}A+D\widehat{B}F=45°+45°=90°\\"

Since two of the angles of quadrilateral DBEF "(E\widehat{D}F \textsf{ and } E\widehat{B}F)" are right angles. DBEF is a square.